OpenStudy (anonymous):

the number N of state and federal inmates in millions during year x, where x ≥ 2002 can be approximated by the following formula. N=0.05x - 99.25 determine the year in which there were 1.1 million inmates.

jimthompson5910 (jim_thompson5910):

you want to know the year when there are 1.1 million inmates so you want to know the value of x (the year) when N = 1.1 (N is number of inmates in millions) so, N=0.05x - 99.25 1.1 = 0.05x - 99.25 .. plug in N = 1.1 now solve for x

OpenStudy (anonymous):

how do I solve for x

jimthompson5910 (jim_thompson5910):

first you want to move that 99.25 over, how do you do that?

jimthompson5910 (jim_thompson5910):

how do you undo the subtraction of 99.25?

OpenStudy (anonymous):

99.25 + 0.05x?

jimthompson5910 (jim_thompson5910):

what do you need to add to both sides

OpenStudy (anonymous):

oh ok so 1.1 + 99.25 = 0.05x - 99.25 + 99.25

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

notice how the - 99.25 + 99.25 on the right side adds up to 0, and it goes away

jimthompson5910 (jim_thompson5910):

so this is effectively like you're "moving" the 99.25 over to the other side

OpenStudy (anonymous):

correct so now i have 1.1 + 99.25 = 0.05x

jimthompson5910 (jim_thompson5910):

now combine the terms on the left side

OpenStudy (anonymous):

which becomes 100.35 = 0.05x

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

then I would divide 100.35 by 0.05 right

jimthompson5910 (jim_thompson5910):

correct, you want to undo the multiplication of 0.05

jimthompson5910 (jim_thompson5910):

so you divide both sides by 0.05

OpenStudy (anonymous):

which would give me the 2007

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

x = 2007

OpenStudy (anonymous):

ok thank you so much seems simple now

jimthompson5910 (jim_thompson5910):

you're welcome, yeah it's not too bad once you get a handle on it and practice it

OpenStudy (anonymous):

I have so many more problems that I dont understand

jimthompson5910 (jim_thompson5910):

well don't hesitate to ask, but keep in mind that you should try some yourself and see how far you get

jimthompson5910 (jim_thompson5910):

ideally there should be a balance between asking questions and doing the work on your own ie, you shouldn't go crazy asking questions, but at the same time, you shouldn't be afraid to ask

OpenStudy (anonymous):

I agree I think I have done good so far on my own

jimthompson5910 (jim_thompson5910):

that's good, you don't want to be completely reliant on others

OpenStudy (anonymous):

I have a total of 200 questions that I am doing and only have 26 left that I havent been able to figure out on my own

jimthompson5910 (jim_thompson5910):

wow that's a lot

OpenStudy (anonymous):

yeah I know

jimthompson5910 (jim_thompson5910):

are the 26 you've yet to figure out related at all?

jimthompson5910 (jim_thompson5910):

or are some related

OpenStudy (anonymous):

only one is related to what we just did

OpenStudy (anonymous):

the rest are different

jimthompson5910 (jim_thompson5910):

ok try to do that one and see if you get anywhere with it

jimthompson5910 (jim_thompson5910):

i gotcha

OpenStudy (anonymous):

I will thanks again

jimthompson5910 (jim_thompson5910):

sure thing